# EquiTabPFN: A Target-Permutation Equivariant Prior Fitted Networks

### v1

@@ Abstract
Recent foundational models for tabular data, such as TabPFN, have demonstrated remarkable effectiveness in adapting to new tasks through in-context learning. However, these models overlook a crucial equivariance property: the arbitrary ordering of target dimensions should not influence model predictions. In this study, we identify this oversight as a source of incompressible error, termed the equivariance gap, which introduces instability in predictions. To mitigate these issues, we propose a novel model designed to preserve equivariance across output dimensions. Our experimental results indicate that our proposed model not only addresses these pitfalls effectively but also achieves competitive benchmark performance.
Machine Learning, ICML
@@ 1 Introduction
Tabular data, a prevalent format in many real-world applications, has historically presented unique challenges to foundational models due to its lack of inherent structure compared to image or text data. Foundation models like TabPFN (Hollmann et al., 2023) have recently emerged, leveraging transformers to address classification problems in tabular settings. These models stand out by eliminating the need for hyperparameter tuning by performing both training and prediction in a single forward pass through a transformer.
The core of those foundational tabular models is a pre-training procedure that trains a transformer to predict on randomly sampled test data given the corresponding (random) training data. Such approaches leverage transformers to perform attention over rows of data, considering all training points simultaneously. Applying attention over rows is critical as it allows us to respect one symmetry of tabular data, e.g. that the order of rows does not matter, which is the case when data is drawn from an i.i.d. distribution (a frequent assumption in supervised learning).
However, row-order symmetry is not the only symmetry relevant to tabular data. Another key symmetry pertains to feature order, where the arrangement of columns should not influence model predictions. Recent work (Müller et al., 2024; Hollmann et al., 2025) has addressed this challenge by employing bi-attention mechanisms similar to those studied in earlier work (Kossen et al., 2022). This approach alternates attention over rows and columns, making the models equivariant to feature permutations and better suited for handling another inherent symmetry of tabular data.
A third important symmetry is the target equivariance of the model: the order in which we encode the target component should not affect the predictions. To the best of our knowledge, target equivariance has not been considered in the context of foundational tabular models. In this paper, we study this aspect and make the following contributions:
  • We demonstrate the critical role of target equivariance in tabular foundational models.
  • We provide a theoretical analysis, showing that the error of TabPFN models can be decomposed into two components: one that measures the model’s equivariance gap and another that evaluates its fit to the data.
  • We propose a novel model that incorporates target equivariance, ensuring more robust and consistent predictions.
  • We show how the architecture we propose improves current foundational tabular architecture in both artificial and real-world datasets.
@@ 2 Related work
Initial work applying transformers to tabular data considered fitting a single transformer per dataset, with attention applied to every column of a single datapoint (Gorishniy et al., 2023). Hollmann et al. (2023) proposed instead to apply attention to the entire dataset at once (e.g., applying attention over the rows of the dataset) and demonstrated how such a model could be trained on synthetic data. This approach has the advantage of requiring only a single forward pass through a transformer to perform both training and predictions. Several works have leveraged this architecture for applications such as Bayesian Optimization (Müller et al., 2023), forecasting (Dooley et al., 2024), learning curve extrapolation (Adriaensen et al., 2023), and fairness analysis (Robertson et al., 2024).
To improve the original architecture, follow-up work proposed new approaches, such as learning the (fast) weights of an MLP capable of performing well on the test portion of a particular dataset (Müller et al., 2023) while achieving much lower latency. Other research focused on modifying the original architecture to align with the natural symmetries of tabular data. For instance, Hollmann et al. (2025); Müller et al. (2024) applied a bi-attention mechanism (Kossen et al., 2022), enabling predictions invariant to the order of features. In particular, the recent work TabPFNv2 (Hollmann et al., 2025) introduced several improvements to the original TabPFN architecture and prior. In addition to being concurrent work, the focus is orthogonal to ours, as we specifically analyze and enhance the output representation of TabPFN.
To the best of our knowledge, no previous approach has proposed a target-equivariant architecture for foundational tabular models, though several works have proposed modifications to the output architecture of TabPFN. Müller et al. (2024) introduced a modification of the output, replacing the linear projection from token embeddings to the target with Generalized Additive Models. This approach improves the interpretability of the model by constructing shape functions for each feature, which can be analyzed independently. Margeloiu et al. (2024) explored training classifiers on features generated by TabPFN. While their method includes non-parametric classifiers such as KNN, the proposed model is not target-equivariant. It also requires training a model at inference time unlike our approach that learns this adaptation during the pre-training.
We are not aware of many theoretical studies analyzing TabPFN, with the exception of (Nagler, 2023). Our work demonstrates the existence of an incompressible term in the pre-training loss, which quantifies the lack of equivariance. Additionally, our architecture enables tabular foundational models to be viewed as (deep) kernel machines, as the output is non-parametric. We believe this perspective opens exciting research directions, given that kernel approaches naturally lend themselves to theoretical analysis (Jacot et al., 2018).
Figure 1: Prediction for both TabPFN and our model when presenting the same datasets but with different class ordering. We pass 9 training points (indicated with dark crosses) each having a distinct class and show the prediction of both models when using 3 different class numbering on a dense grid. The predictions of our model are identical under different class numbering as opposed to TabPFN as the model is not equivariant.
@@ 3 Background on Prior-Fitted Networks
Hollmann et al. (2023) introduced a pre-trained model, TabPFN, that leverages the transformer architecture (Vaswani et al., 2017) to perform in-context learning on unseen tabular datasets without the need for any further training. Below, we briefly review the three main ingredients for pre-training of this model as it will help to describe the architectural changes we propose to make the model equivariant to target permutations.
3.1 Dataset prior
A key ingredient to the pre-training procedure is the use of artificial datasets that are sampled using a sophisticated generative model meant to capture the real-world distribution of datasets. More precisely, any dataset of multiple covariate/target pairs $(x,y)$ is generated by first sampling a latent parameter $\psi$ characterizing the dataset according to a pre-defined prior $p(\psi)$, then sampling each covariate/target pair $(x,y)$ in the dataset according to a conditional distribution $p(x,y|\psi)$ characterized by $\psi$. The pre-training procedure requires repeatedly generating two collections of training and test samples $(x_{n},y_{n})_{n=1}^{N}$ and $(x_{m}^{\star},y_{m}^{\star})_{m=1}^{M}$ all of which are sampled from the conditional distribution $p(x,y|\psi)$ using a shared value of $\psi$ sampled from the prior $p(\psi)$. Here, we will use the same generative mechanism as Hollmann et al. (2023) for pre-training the proposed equivariant model.
@@ 3.2 Pre-training procedure
TabPFN is trained by predicting test target values $(y_{m}^{\star})_{m=1}^{M}$ given corresponding test covariates $(x_{m}^{\star})_{m=1}^{M}$ as well as training covariates/target pairs $(x_{n},y_{n})_{n=1}^{N}$. Denoting for simplicity $(X,Y)$ and $(X^{\star},Y^{\star})$, the collection of training and test samples, TabPFN’s pre-training objective is given by:
$\mathcal{L}(f):=\mathbb{E}\left[\ell\left(f_{X,Y}\left(X^{\star}\right),Y^{\star}\right)\right],$ (1)
@@
where $(y,y^{\prime})\mapsto\ell(y,y^{\prime})\in\mathbb{R}$ is a point-wise loss, while $f_{X,Y}\left(X^{\star}\right)$ is the output of the network when provided with the training collection $(X,Y)$ and test queries $X^{\star}$. Here, the expectation is over the collections $(X,Y)$ and $(X^{\star},Y^{\star})$ sampled according to the dataset prior. The loss $\ell$ is, typically, chosen to be the mean-squared-error loss in the case of regression or the cross-entropy loss in the case of classification. Note that each test query $x_{m}^{\star}$ is processed independently by the network $f$ so that $f_{X,Y}\left(X^{\star}\right)=(f_{X,Y}\left(x_{m}^{\star}\right))_{m=1}^{M}$. The above procedure trains the model to directly infer new covariate/target dependences from the provided training pairs and will be used in our setting as well.
3.3 TabPFN architecture
The architecture used in TabPFN can be decomposed as into three modules, an encoder, a backbone and a decoder.
Linear Encoder. The encoder module constructs training and test tokens $(e_{n})_{n=1}^{N}$ and $(e^{\star}_{m})_{m=1}^{M}$ that are provided to the transformer backbone assuming the inputs $x$ and $y$ are vectors of fixed dimensions $p$ and $q_{max}$. Each training token $e_{n}$ is obtained by linearly embedding both covariate $x_{n}$ and target $y_{n}$ into a feature space of fixed dimension $d$ and then summing both embeddings, i.e. $e_{n}=Ux_{n}+Vy_{n}$, where $U$ and $V$ are trainable matrices of sizes $d\times p$ and $d\times q$. On the other hand, the test token consists only in embedding the test covariate $x^{\star}_{m}$, i.e. $e^{\star}_{m}=Ux_{m}^{\star}$ since the test target $y^{\star}_{m}$ is not provided to the network. While targets with smaller dimensions can be handled by a simple zero-padding procedure (see Hollmann et al. (2023)), the encoder cannot easily accommodate target data with dimensions greater than $q$.
MLP decoder. The decoder consists of a one-hidden layer MLP that takes each test output token $e^{\star}_{m}$ produced by the transformer backbone and produces a prediction vector $\hat{y}_{m}$ of fixed dimension $q$. This choice of decoder requires a pre-defined target dimension $q$. Thus, it cannot be used post-hoc on new data with higher target dimensions.
A notable property of the TabPFN architecture is its invariance of the test predictions to the order by which training and test points are provided. This property is desirable since the order of training points is arbitrary and should not influence predictions on test samples. However, the network lacks equivariance w.r.t. the targets’ dimensions, meaning that predictions are highly dependent on the order by which the training target dimensions are provided, an undesirable property as we show next.
4 Target permutation equivariance in PFN
4.1 Target permutation equivariance

### v2

@@ Abstract
Recent foundational models for tabular data, such as TabPFN, excel at adapting to new tasks via in-context learning, but remain constrained to a fixed, pre-defined number of target dimensions—often necessitating costly ensembling strategies. We trace this constraint to a deeper architectural shortcoming: these models lack target equivariance, so that permuting target dimension orderings alters their predictions. This deficiency gives rise to an irreducible “equivariance gap,” an error term that introduces instability in predictions. We eliminate this gap by designing a fully target-equivariant architecture—ensuring permutation invariance via equivariant encoders, decoders, and a bi-attention mechanism. Empirical evaluation on standard classification benchmarks shows that, on datasets with more classes than those seen during pre-training, our model matches or surpasses existing methods while incurring lower computational overhead.
@@ 1 Introduction
Tabular data, a prevalent format in many real-world applications, has historically presented unique challenges to deep learning due to its lack of inherent structure compared to image or text data [7]. Foundation models, such as TabPFN [9], have recently been introduced to tackle classification tasks in tabular domains. These models leverage in-context learning capabilities of transformers [3], to perform both training and prediction in a single model evaluation, without requiring any parameter updates, achieving remarkable performance.
At the core of these foundational models is a pre-training procedure in which a transformer model is trained to predict test targets from test covariates, conditioned on training covariate/target pairs all of which are randomly sampled from some well-designed generative model. While it might seem surprising, at first, how a model pre-trained on synthetic data could perform well on real unseen data, recent work, such as Nagler [21], provides a theoretical study that shades some light on this phenomenon. These models, leverage the transformer architecture to perform attention over rows—attending to all samples simultaneously to enable cross-sample comparison. Applying attention over rows is crucial, as it enables the model to capture higher-order similarities between samples while preserving an inherent symmetry of tabular data—namely, that row order is irrelevant under the common i.i.d. assumption in supervised learning. Conveniently, such a mechanism also allows, in theory, handling an arbitrary number of training samples—a property that enables the model to generalize across tasks with varying dataset sizes without architectural modifications.
However, models such as TabPFN [9] are inherently confined to covariate–target pairs of fixed, predefined dimension, thereby limiting their applicability to datasets that match these specifications. This limitation can be alleviated via additional data pre-processing—e.g., projecting high-dimensional covariates into a lower-dimensional subspace—or by post-processing model predictions,—e.g. employing hierarchical strategies for classification tasks with many classes [24, 4]. However, the increased computational burden often offsets some of the benefits offered by in-context learning.
Recent work in [19, 10] have partially addressed these challenges, allowing the model to handle arbitrary number of covariates, albeit, requiring the target dimension to have a fixed pre-defined dimension. A key insight there, is to exploit another inherent symmetry of tabular data: the arrangement of columns/covariates’s dimensions should not influence model predictions. This is achieved through the bi-attention mechanism which alternates between attention over samples/rows and columns, i.e. covariates dimensions, thereby making the model equivariant to feature permutations just as it is equivariant to sample permutations. Nevertheless, these models remain limited to tasks where the target size matches the predefined dimensionality. While covariates are provided for both training and test samples, only training targets are available to the model. This asymmetry between training and test samples complicates direct extensions of the above approaches to handle the targets.
In this work, we propose EquiTabPFN, a novel architecture that enforces target equivariance, thus ensuring more robust and consistent predictions while handling targets of arbitrary dimensions. Unlike feature equivariance which can be directly obtained using a bi-attention mechanism, we achieve target equivariance by carefully combining three different mechanisms: (1) Bi-attention across covariates/target components and datapoints, (2) Prediction tokens to replace unavailable test targets, and (3) A non-parametric decoder preserving equivariance in predictions. We then establish the importance of target equivariance through theoretical and empirical analyses. In our theoretical study, we show that optimal functions for the pre-training procedure must necessarily be target-equivariant. Finally, we demonstrate, on real-world datasets, that target-equivariant models are beneficial both in terms of classification performance and inference runtime.
@@ 2 Related Work
Prior-Fitted Networks. Since the introduction of TabPFN in Hollmann et al. [9], which demonstrates how such a model can be successfully trained on synthetic data, several works have leveraged this architecture for applications such as Bayesian Optimization [18], forecasting [6], learning curve extrapolation [1], and fairness analysis [23]. Aside from Hollmann et al. [10], Müller et al. [19] that proposed a covariate equivariant version of the TabPFN models to better capture natural symmetries of tabular data, other works focused on improving scalability and speed of the model. This includes Müller et al. [17] who proposed to pre-train the model for producing the weights on an MLP by in-context learning, so that the resulting MLP performs well on a test portion of a particular dataset while achieving much lower latency. Recently, Qu et al. [22] improved the scalability w.r.t. to the sample size by introducing a two-stage architecture that first builds fixed-dimensional embeddings of rows, followed by a transformer for efficient in-context learning. All these approaches, still require a pre-defined fixed target dimension. The focus is orthogonal to ours, as we specifically analyze and enhance the target’s representation of TabPFN family of architectures.
Beyond pre-defined target dimensionality. Margeloiu et al. [14] explored training classifiers on features generated by TabPFN. This requires training a model at inference ; unlike our approach. To address TabPFN’s limitation to a predefined number of classes, Hollmann et al. [10] propose to split the classification problem into smaller ones on which the model can be employed, then to aggregate predictions using a strategy, such as the one based on an error-correcting output codes (ECOC) [4]. In this work, we show that such a strategy results in an increased computational cost compared to using our architecture that is natively target equivariant. Qu et al. [22] proposed to use a hierarchical classification strategy [24] which still incurs an increased computational cost. Recently, Wu and Bergman [28] propose a mechanism to handle arbitrary number of classes without target equivariance and require a different paradigm involving an adversarial training procedure.
@@ 3 Background on Prior-Fitted Networks
Hollmann et al. [9] introduced a pre-trained model, TabPFN, that leverages the transformer architecture [26] to perform in-context learning on unseen tabular datasets for classification tasks without the need for any further training. Specifically, given training and test datasets of input-output pairs $(X,Y):=(x_{n},y_{n})_{n=1}^{N}$ and test samples $(X^{\star},Y^{\star}):=(x_{m}^{\star},y^{\star})_{m=1}^{M}$, TabPFN returns a prediction $\hat{Y}=f_{X,Y}\left(X^{\star}\right)$, where $f_{X,Y}\left(X^{\star}\right)$ is the output of the network when provided with the training collection $(X,Y)$ and test queries $X^{\star}$. Here the input vectors $x_{n}$ and $x_{m}^{\star}$ belong to a euclidean space $\mathbb{R}^{p}$, while classes are represented by one-hot vectors $y_{n}\in\mathbb{R}^{q}$.
We now briefly describe the three modules of TabPFN model: an encoder, a backbone and a decoder, as it will help identify the main architectural constrains that impose a limit on the number of classes.
Linear Encoder. The encoder module constructs training and test tokens $(e_{n})_{n=1}^{N}$ and $(e^{\star}_{m})_{m=1}^{M}$ that are provided to the transformer backbone assuming the inputs $x$ and $y$ are vectors of fixed dimensions $p$ and $q_{max}$. Each training token $e_{n}$ is obtained by linearly embedding both covariate $x_{n}$ and target $y_{n}$ into a feature space of fixed dimension $d$ and then summing both embeddings, i.e. $e_{n}=Ux_{n}+Vy_{n}$, where $U$ and $V$ are trainable matrices of sizes $d\times p$ and $d\times q$. On the other hand, the test token consists only in embedding the test covariate $x^{\star}_{m}$, i.e. $e^{\star}_{m}=Ux_{m}^{\star}$ since the test target $y^{\star}_{m}$ is not provided to the network. While targets with smaller dimensions can be handled by a simple zero-padding procedure (see Hollmann et al. [9]), the encoder cannot easily accommodate target data with dimensions greater than $q$.
MLP decoder. The decoder consists of a one-hidden layer MLP that takes each test output token $e^{\star}_{m}$ produced by the transformer backbone and produces a prediction vector $\hat{y}_{m}$ of dimension $q$. As the decoder requires a pre-defined target dimension $q$, it cannot be used post-hoc on new data with higher target dimensions.
A notable property of the TabPFN architecture is its invariance of the test predictions to the order by which training and test points are provided. This property is desirable since the order of training points is arbitrary and should not influence predictions on test samples. However, the network lacks equivariance w.r.t. the targets’ dimensions, meaning that predictions are highly dependent on the order by which the training target dimensions are provided, an undesirable property as we show in our theoretical analysis in Section 5.
4 Target equivariant prior-fitted network
Figure 1: Overview of EquiTabPFN’s architecture. Data is tokenized via an encoder, processed using self-attention, and decoded to obtain predictions. The encoder maps each covariate to a single token and embeds target components into tokens via a $1\times 1$ convolution. Missing test tokens are replaced by prediction tokens. Self-attention alternates between (1) feature-wise attention, with target tokens attending only to covariate tokens (gray arrows) while covariate tokens attend to all tokens (blue arrows); (2) Data-wise attention, where test tokens attend only to training tokens (blue arrows), and training tokens attend to themselves (gray arrows).
We introduce EquiTabPFN, a new model architecture for in-context learning on tabular data, that is permutation equivariant w.r.t. the target’s components. Our architecture integrates self-attention mechanisms across data points and data components to leverage relationships between datapoints while preserving equivariance by processing individual attributes (such as the targets components). Unlike TabPFN which requires fixed target dimensions for all datasets, EquiTabPFN allows the dimensions of the target to change depending on the dataset as a consequence of its equivariant architecture. EquiTabPFN consists of three major modules: a target equivariant encoder, an attention module both across data components and across data points, and a non-parametric equivariant decoder, each designed to facilitate the end-to-end learning of components interactions and datapoint relationships, see Figure 1. Below, we elaborate on each module and their interactions.
4.1 Target equivariant encoder
The encoder constructs training and test tokens by applying a linear projection to both covariates and targets so that target equivariance is preserved. Specifically, following [9], each training and test covariate vector $x_{n}$ and $x^{\star}_{m}$ is encoded into a single token of dimension $d$ by applying a linear projection matrix $U$ of size $d\times p$. However, instead of adding a linear projection of the training targets to each corresponding training token, as done in the case of TabPFN (see Section 3), we compute a token for each component $(y_{n})_{j}$ of a training target by multiplying them with an embedding vector $V$ of dimension $d$ for all $1\leq j\leq q$. This operation amounts to applying a $1\times 1$ convolution along the components of each target which preserves target equivariance. Since, the validation target $Y^{\star}$ is not provided to the model as input, it is replaced by a trainable prediction token $W_{\text{pred}}$ of dimension $d$ that is repeated $M\times q$ times to form an initial guess $\tilde{Y}^{0}$ of the target. All these embeddings along with prediction tokens are collected to form a single tensor $E$ of shape $(N+M,q+1,d)$, where the blocks $E_{:N,1,:}$, $E_{N:M,1,:}$ correspond to embeddings of $X$ and $X^{\star}$, $E_{:N,1:q,:}$ represents the embedding of $Y$ while $E_{N:M,1:q,:}$ denotes the initial guess $\tilde{Y}^{0}$ obtained using the prediction token. This tensor is then processed by the attention modules as described next. inline]DS: Write matrix dimensions
4.2 Self-Attention Mechanisms
The core of the architecture involves two alternating self-attention modules: self-attention across components $\text{SelfAtt}_{c}$ and self-attention across datapoints $\text{SelfAtt}_{b}$ used for transforming the tokens. These alternating self-attention layers allow the model to learn both intra-samples components interactions and inter-samples relationships. Following standard design choices for transformers, we apply residual connections and layer normalization to ensure stability and robust gradient flow, i.e.:
$E$ $\leftarrow\text{LN}\left(E+\textbf{SelfAtt}_{c/b}(E)\right),$ $E$ $\leftarrow\text{LN}\left(E+\text{MLP}(E)\right),$
where LN denotes the layer normalization layer [2], $\text{SelfAtt}_{c/b}$ denotes one of the considered self-attention mechanisms and MLP is a one hidden-layer network acting on each embedding independently. Below, we describe both self-attention mechanisms in more detail.
Self-attention across components
Self-Attention across datapoints
captures relationships between datapoint embeddings, allowing the model to aggregate information globally. It is applied between samples and independently per each input dimensions $p$ and $q$ to preserve equivariance. Similarly to [9], training and validation tokens only attend to training tokens.
4.3 Non-parametric equivariant decoder
The decoder aggregates the processed embeddings to produce prediction $\hat{Y}$. This is achieved in two steps: an attention module first computes an intermediate prediction $\tilde{Y}=(\tilde{y}_{m})_{m=1}^{M}$ in the form of a weighted average of training targets $Y$, then a residual correction is added to produce the final prediction. More precisely, the attention module uses the embeddings of the training and validation samples as keys and queries, while the attention values are simply the training targets $Y$, i.e. $\tilde{y}_{m}=\sum_{n=1}^{N}y_{n}\text{SoftMax}\left(\sum_{i,u}E_{n,i,u}E_{m,i,u}/\sqrt{(1+q)d}\right).$ The residual correction, in the form of a point-wise MLP, operates independently on each dimension $j$ of the attention output $\tilde{y}_{m}$ so that equivariance is preserved while enabling nonlinear interactions between training values.
Without the residual correction and removing the dependence of the keys and queries embeddings on the training targets $Y$ (for instance by setting the weights of the target encoder and pointwise MLP to 00), the decoder becomes a linear non-parametric regression estimator [25, Definition 1.7], which is a generalization of Nadaraya-Watson’s estimator [20, 27]. However, linear estimators are known to be suboptimal compared to non-linear ones [5]. This motivates introducing a nonlinear dependence of the estimator to $Y$, in our setting, to increase the expressiveness of the decoder allowing it to adapt to the prediction task at hand. Experimentally, we also found a clear improvement when adding such a residual correction and making the embeddings dependent on the targets.
@@ 4.4 Pre-training Procedure
EquiTabPFN can be pre-trained using the same procedure as in Hollmann et al. [9], on artificial datasets sampled from a sophisticated generative model meant to capture the real-world distribution of datasets. More precisely, each artificial dataset consists of training/and test splits $(X,Y):=(x_{n},y_{n})_{n=1}^{N}$ and $(X^{\star},Y^{\star}):=(x_{m}^{\star},y_{m}^{\star})_{m=1}^{M}$ sampled according to a conditional distribution $p(x,y|\psi)$ characterized by a latent parameter $\psi$. The parameter $\psi$ characterizes the dataset and is itself sampled according to a predefined prior $p(\psi)$. The pre-training procedure requires repeatedly generating artificial datasets and training the model to predict test target values $Y^{\star}$ given corresponding test covariates $X^{\star}$ as well as training covariates/target pairs $(X,Y)$ by minimizing an objective of the form:
$\mathcal{L}(f):=\mathbb{E}\left[\ell\left(f_{X,Y}\left(X^{\star}\right),Y^{\star}\right)\right].$ (1)
@@
Here, $(y,y^{\prime})\mapsto\ell(y,y^{\prime})\in\mathbb{R}$ is a point-wise loss, typically cross-entropy, and the expectation is over the collections datasets sampled according to the dataset prior. Note that each test query $x_{m}^{\star}$ is processed independently by the network $f$ so that $f_{X,Y}\left(X^{\star}\right)=(f_{X,Y}\left(x_{m}^{\star}\right))_{m=1}^{M}$. Next, we show, under natural conditions, that target equivariant functions constitute the right space of functions when searching for solutions to objectives of the form in Equation 1.
5 Target permutation equivariance and prior-fitted networks
We now formally analyze the impact of non-target equivariance on the training objective. We begin by precisely defining target equivariance, then show that an optimal solution must be equivariant or otherwise incurs an error quantified by the target equivariance gap. Finally, through empirical analysis of TabPFN training, we illustrate how the equivariance gap decreases slowly, highlighting the fundamental challenge non-equivariant architectures face in learning this key data symmetry.
5.1 Optimality of target equivariant networks